# RU-AI: A Large Multimodal Dataset for Machine Generated Content Detection

### v1

@@ Abstract.
The recent advancements in generative AI models, which can create realistic and human-like content, are significantly transforming how people communicate, create, and work. While the appropriate use of generative AI models can benefit the society, their misuse poses significant threats to data reliability and authentication. However, due to a lack of aligned multimodal datasets, effective and robust methods for detecting machine-generated content are still in the early stages of development. In this paper, we introduce RU-AI, a new large-scale multimodal dataset designed for the robust and efficient detection of machine-generated content in text, image, and voice. Our dataset is constructed from three large publicly available datasets: Flickr8K, COCO, and Places205, by combining the original datasets and their corresponding machine-generated pairs. Additionally, experimental results show that our proposed unified model, which incorporates a multimodal embedding module with a multilayer perceptron network, can effectively determine the origin of the data (i.e., original data samples or machine-generated ones) from RU-AI. However, future work is still required to address the remaining challenges posed by RU-AI. The source code and dataset are available at https://github.com/ZhihaoZhang97/RU-AI.
@@
ccs: Information systems Multimedia content creation
@@ 1. Introduction
The recent advent of large generative AI models has significantly impacted various sectors of human society, including industry (Cardon et al., 2023; Dwivedi et al., 2023; Epstein et al., 2023), research (Walters and Murcko, 2020; Megahed et al., 2023; Peres et al., 2023), and education (Qadir, 2023; Peres et al., 2023). Current generative AI models are capable of producing realistic images (e.g. Stable Diffusion (Rombach et al., 2022), MidjourneyV5 (Holz et al., [n. d.]), DALL-E2 (Ramesh et al., 2022)), human-level text (e.g. ChatGPT (Ouyang et al., 2022), GPT-4 (OpenAI, 2023), LLaMa (Touvron et al., 2023a)) and natural-sound voice (e.g. YourTTS (Casanova et al., 2022), VITS (Kim et al., 2021)), which covers the major multimedia information in communication, creation and work. Deploying generative AI models appropriately can save professionals’ time for substantive matters (Hacker et al., 2023), assist scientists in discovering new medicine or compounds (Zhavoronkov et al., 2019), and provide students with customised explanations and feedback (Grassini, 2023). However, generative AI models’ capabilities may be misused for generating discriminate or disrespectful content (Borgesius, 2020), spreading misinformation (Zhou et al., 2023), and conducting target scams (Wang et al., 2020a). Hence, it is important to identify whether a piece of content is from a human or a machine.
@@
Previous studies have explored manual methods and automated approaches for detecting machine-generated content. Researchers (Mori et al., 2019; Samo and Highhouse, 2023; AlBadawy et al., 2019) have found that humans can identify machine-generated content with substantial accuracy, but the quality of results is largely influenced by an individual’s background, experience and preference. Moreover, human experts may experience difficulties in distinguishing machine-generated content from real content in specific domains, such as creative writing (Gunser et al., 2021), art painting (Elgammal et al., 2017) and unfamiliar voices (Mai et al., 2023). In addition, identifying machine-generated content manually is labour-intensive and challenging to scale, which results in machine-generated content not being identified promptly. Therefore, it is important to explore automated approaches capable of detecting machine-generated content across different modalities, with a balance between accuracy and efficiency. In practice, corresponding comprehensive datasets form the foundation for developing automated detection approaches and models. However, the majority of the existing datasets for machine-generated content detection (Fagni et al., 2021; Wang et al., 2023; Liu et al., 2023; Li et al., 2023b; Guo et al., 2023; Wang et al., 2020a; Frank and Schönherr, 2021; Salvi et al., 2023; Göring et al., 2023; Wang et al., 2020b; Zhu et al., 2023) are constructed within a single modality, disregarding valuable information across modalities that could aid in detecting the machine-generated content. Although there are several multimodal datasets for machine-generated content detection (Khalid et al., 2021; Dolhansky et al., 2019; He et al., 2021), they mainly contain information on human faces and voices, significantly limiting the scope of their usage. Only one dataset (Shao et al., 2023) contains machine-generated images and text in the general domain, but it still lacks of machine synthesised voices.
To address these problems, we introduce RU-AI, a large-scale multimodal dataset for robust detection of machine-generated content in text, image and voice, and develop baseline detection models and perform extensive experiments with them. The real portion of our multimodal dataset includes three large publicly available image-caption datasets: COCO (Lin et al., 2014), Flickr8K (Young et al., 2014) and Places205 (Zhou et al., 2014), along with their corresponding human voice transcriptions. Machine-generated data is produced by inferencing generative AI models, including five models for text generation, image generation, and voice generation. This yields 245,895 aligned instances each for both real and machine-generated data. Moreover, to prove the effectiveness of our dataset, we develop and train baseline detection models on RU-AI. Our contribution can be summarised as follows:
  • We introduce a newly constructed large-scale multimodal dataset that comprises true data and machine-generated data, with instances aligned across text, image, and voice modalities. The dataset contains multiple categories of objects and the machine-generated data is produced by using five generative AI models for each modality.
  • We develop baseline detection models and conduct experiments by testing the performance of them across all modalities, demonstrating the effectiveness and values of our constructed datasets in training and testing models for machine-generated content detection.
@@ 2. RU-AI Dataset
We introduce RU-AI, a new large-scale dataset that features multimodal real and human-generated content, accompanied by machine-generated content. It encompasses three modalities: real-world images, image-describing caption text, and spoken voices for captions, providing more than 245k real and machine-generated aligned instance pairs for each modality. More specifically, image and text instances are collected from COCO (Lin et al., 2014), Flickr8K (Young et al., 2014) and Places205 (Zhou et al., 2014), while voice instances are sourced from SLS Corpora (Harwath et al., 2016; Harwath and Glass, 2017; Hsu et al., 2021; Harwath and Glass, 2015), which form the real and human-generated part of our dataset. The machine-generated portion is produced by five state-of-the-art generative models specific to each modality. Figure 1 presents the data construction pipeline.
@@ 2.1. Data Construction
Voice. The voices for captions are cloud-sourced from a group of individuals, with each individual being assigned random captions (Harwath et al., 2016; Harwath and Glass, 2017; Hsu et al., 2021; Harwath and Glass, 2015). Some text-to-speech (TTS) models require reference voices for inference; therefore, we merge the spoken caption voices of each individual, identified by their unique user ID, into one-minute long voice files. For YourTTS (Casanova et al., 2022), the machine-generated voices are synthesised by using the captions along with their corresponding one-minute human voices as references. For other models, the voices are synthesised directly from the caption text using default settings. All five models contribute equally to the generated voice data, and these voices are named and saved similarly to the images, with the original IDs and model names attached.
@@ 2.2. Data Processing
LLMs may add extra information such as ”Sure, here is the sentence:” and ”Certainly, here’s a rephrased version of your text:” during the rephrasing process, which is unrelated to the generated sentences. We utilize a regular expression method to remove this information from the generated sentences to ensure the quality of the text data. Since censorship can be applied to generative models for safety and alignment, empty sentences or blank images can appear during the generation process. We remove them to ensure consistency across our dataset. To ensure alignment across the three modalities, we perform a data-matching process shown in Algorithm 1, where real and human-generated content is represented by $R$, and machine-generated content is represented by $N$. The generated images and captions, represented by $N_{image}$ and $N_{caption}$, are a subset of the real images and captions, while $R_{voice}$ corresponds to human-spoken voices, having the same amount of data as the human-generated ones. For each instance, we retrieve the unique IDs of the real images and captions ($R_{image_{i}}$ and $R_{caption_{i}}$) and verify whether they both exist in the generated sets. Since we intentionally ensure that the original and machine-generated data share the same unique ID, a matched data instance is formed by aligning the image, text, and audio that correspond to the specific ID. We implement this approach in our collected dataset, resulting in 245,895 aligned real-and-AI data pairs. As shown in Table 1, the final processed data for each modality includes 7,890 pairs from Flickr8K, 119,264 pairs from COCO, and 118,741 pairs from Places205.
Algorithm 1 Data matching across text, image and voice
$R_{caption},R_{image},R_{voice}$
$N_{image},N_{caption},N_{voice}$
while $N_{image}$, $N_{caption}$ do
     if $N_{image_{i}}\in R_{image}\And N_{caption}\in R_{caption}$  then
         $M_{match_{j}}\leftarrow(N_{image_{i}},R_{image})\And$
         $M_{match_{j}}\leftarrow(N_{caption_{i}},R_{caption_{i}})\And$
         $M_{match_{j}}\leftarrow(N_{voice_{i}},R_{voice_{i}})$
     end if
end while
return $M_{match}$
@@
Text ChatGPT (Ouyang et al., 2022) 1,610 24,272 23,736
Llama-2-13B (Touvron et al., 2023b) 1,518 23,121 23,488
@@
Mixtral-8x7B (Jiang et al., 2024) 1,599 23,752 23,858
PaLM-2 (Anil et al., 2023) 1,535 24,221 23,879
@@
AbsoluteReality-1.8  [^1] Both authors contributed equally to this research. 1,628 23,898 23,780
@@
StableDiffusion-XL (Podell et al., 2023) 1,535 24,221 23,879
Audio YourTTS (Casanova et al., 2022) 1,610 24,272 23,736
XTTS-2(Golge and Team, 2021) 1,518 23,121 23,488
@@ 2.3. Data Characteristics
Our dataset includes 245,895 pairs of real and machine-generated data from text, image, and voice modalities, resulting in a total of 1,475,370 instances collected from three diverse datasets. Table 2 shows the average length of captions from different LLMs and their average similarity between the reference and generated text, indicated by average ROUGE-L and BERTScore. Table 3 shows the synthesised voice characteristics from different models, including average voice duration, average amplitude, and average spectral flatness. The average amplitude range refers to the average normalised audio scale, while the average spectral flatness is a measure of the sound’s noise-like level. Figure 2 displays the object categories featured in images, which also serve as the topics in text and voice modalities. Since all images in Flickr8K are unlabelled with object instances, we employ pre-trained YOLOv5 (et. al., 2021) to annotate the unlabelled images. The data characteristics are based on the original annotations and the YOLOv5 annotations. Our dataset can be divided into 20 different categories, with ”Person” being the largest category, and there are 1,040 images that do not belong to any category and are classified as ”Other”.
Figure 2. Distribution of image data in each category.
@@ 3. Proposed Baseline
Implementation Details. We use two pre-trained multimodal encoder models with their large settings as the baseline multimodal embedding models: ImageBind (Girdhar et al., 2023) and LanguageBind (Zhu et al., 2024). ImageBind encodes all input data into 1,024-dimensional embedding sequences, while LanguageBind encodes them into 768-dimensional embedding sequences. Each multimodal embedding model is connected to a separate multilayer perceptron network, consisting of a single 256-dimensional hidden layer. The input layers of these networks match the dimensions of their respective embedding models, 1,024-dimensional for ImageBind and 768-dimensional for LanguageBind. These networks are used for the classification of the input data as either real or machine-generated one. Specifically, the output is calculated from probability vectors $\vec{z}=(z_{r},z_{m})$ using the softmax function $\sigma(\vec{z})_{i}$. The instance is classified as real if $z_{r}>z_{m}$; otherwise, it is classified as machine-generated. To analyse the effectiveness of different modalities in our data, we train and evaluate our classification model with different combinations of the data. The training samples are randomly selected from 80% of each data combination, while the remaining 20% of the samples are used as testing sets. We maintain the same size of data throughout the experiment and distribute the modalities equally. The performance is evaluated using Accuracy, Precision, and F1 metrics on the trained-modality (TM) and all-modalities (ALL). For the baseline implementation, we use PyTorch 1.13 with CUDA 11.8 on a server with a single Nvidia RTX 3090 GPU. All models were trained for 5 epochs, using the Adam optimizer with cross-entropy loss and a learning rate of 1e-4, with the embedding model frozen. We follow the embedding models’ original feature extraction techniques in all three modalities for consistency across datasets.
Results.The classification results from our proposed model are shown in Table 4. We observe that voice is the simplest modality to distinguish between human and machine-generated content. This aligns with the model behaviours in all-modality evaluation, as the accuracy and precision are the lowest when the model is trained only with voice data. This is likely attributable to the obvious feature difference between human and machine-generated voices. Combining data from different modalities improves the model’s performance in all-modality evaluation but affects its effectiveness on the trained modality. This is expected since mixing modality data is more complex than single modality data, which could cause difficulty in model convergence. Models utilising LanguageBind as embedding models generally perform better than those incorporating ImageBind as embedding models. This might be due to LanguageBind’s enhanced cross-modality architecture, which binds all modalities through language with rich semantics to provide more comprehensive information even with a smaller embedding size. According to our baseline results, machine-generated content detection remains a challenging task, especially in terms of text and cross-modal detection. To subside the potential misuse of generative models, future work is required to explore more robust and reliable methods for detecting machine-generated content.
@@
Table 4. Results of the proposed classification model. ”LB” is LanguageBind, ”IB” is ImageBind, ”A” is Accuracy, ”P” is Precision, ”ALL” is all three modalities, ”TM” is trained-modality.
Model Modality A@TM P@TM F1@TM A@ALL P@ALL F1@ALL
LB + MLP Text 75.04 75.02 75.04 69.76 69.78 69.75
Image 94.74 93.78 94.79 76.49 85.95 72.93
Voice 100.00 100.00 100.00 68.89 64.09 73.43
Text+Image 74.05 73.15 74.55 77.09 76.91 77.16
Text+Voice 86.25 85.57 86.38 77.69 81.56 76.23
Image+Voice 96.97 96.88 96.97 82.26 78.89 83.23
All N/A N/A N/A 84.20 84.20 82.67
IB + MLP Text 80.83 79.91 81.12 60.70 57.36 67.96
Image 95.23 94.66 95.26 67.75 71.03 65.03
Voice 100.00 100.00 100.00 65.77 63.61 68.28
Text+Image 79.37 77.97 79.87 71.48 75.18 69.23
Text+Voice 87.95 87.97 87.95 75.39 69.05 78.90
Image+Voice 96.97 97.02 96.96 81.25 82.84 80.78
All N/A N/A N/A 80.63 79.19 81.10

### v2

# RU-AI: A Large Multimodal Dataset for Machine-Generated Content Detection

@@ Abstract.
The recent generative AI models’ capability of creating realistic and human-like content is significantly transforming the ways in which people communicate, create and work. The appropriate use of generative AI models can benefit society, while their misuse poses threats to the society. However, the lack of aligned multimodal datasets has inhibited the development of effective and robust methods for detecting machine-generated content, particularly in triple-modality settings (e.g., text, image, and voice). In this paper, we introduce RU-AI, a new large-scale multimodal dataset for robust and efficient detection of machine-generated content in text, image and voice. Our dataset is constructed on the basis of three large publicly available datasets: Flickr8K, COCO and Places205, by adding their corresponding AI duplicates, resulting total of 1,475,370 data instances. In addition, we create a noise variant of each modality of the datasets aiming to analyse the models’ robustness. Given our dataset, we conduct extensive experiments with the current SOTA detection methods. The results reveal that existing models still struggle to achieve accurate and robust classification after training on our dataset. The RU-AI dataset is designed to support the development of detection methods across modalities and can be effectively utilised for identifying machine-generated content. The source code and dataset are available at https://github.com/ZhihaoZhang97/RU-AI.
@@
conference: Make sure to enter the correct conference title from your rights confirmation email; April 28–May 02, 2025; Sydney, Australiaccs: Information systems Multimedia content creation
@@ 1. Introduction
The recent advent of large generative AI models has significantly impacted various sectors of human society, including industry, research, and education. They are capable of producing realistic images, human-level text and natural-sound voice, which covers the major multimedia information in communication, creation and work. Deploying generative AI models appropriately can save professionals’ time for substantive matters, assist scientists in discovering new medicine or compounds, and provide students with customised explanations and feedback. However, generative AI models’ capabilities may be misused for generating discriminate or disrespectful content, spreading misinformation, and conducting target scams. These challenges indicate the pressing need for robust mechanisms and models to verify data originality, particularly for multimodal content.
Previous studies  (Jawahar et al., 2020) have explored manual methods and automated approaches for detecting machine-generated content. Researchers (AlBadawy et al., 2019) have found that humans can identify machine-generat-ed content with substantial accuracy, but the quality of results is largely influenced by an individual’s background, experience and preference. Additionally, manually identifying machine-generated content is labour-intensive and difficult to scale, while detection across multiple modalities is inherently more challenging than within a single modality. Therefore, it is curial to explore automated approaches to detect machine-generated content across different modalities. However, the majority of existing datasets for machine-generated content detection (Zhu et al., 2023) are constructed within a single modality, thereby neglecting valuable cross-modal information that could enhance the detection of machine-generated content. Although there exists several multimodal datasets for machine-generated content detection (Khalid et al., 2021), they primarily focus on human faces and voices, which significantly limits their scope of applicability. Only one dataset (Shao et al., 2023) includes machine-generated images and text, but lacks synthesised voices. Combining audio and images with text helps to reduce ambiguity, while integrating all three modalities can minimise information loss. This approach can be effectively applied to combat fraud and disinformation.
@@
To address these problems, we introduce RU-AI, a large-scale multimodal dataset for robust detection of machine-generated content in text, image and voice. The real portion of our multimodal dataset are collected from three large publicly available image-caption datasets: COCO (Lin et al., 2014), Flickr8K (Young et al., 2014) and Places205 (Zhou et al., 2014) with their human voices. Machine-generated data is produced from generative AI models, including five models for text, image, and voice generation. This yields total of 1,475,370 instances each for both real and machine-generated data. We also construct a noise-augmented variant of our dataset across all modalities to evaluate model robustness. Furthermore, to demonstrate the effectiveness of our dataset, we develop and train baseline detection models on RU-AI and conduct extensive experiments to assess their performance on our dataset. Our key contributions can be summarised as follows:
  • We introduce the first large-scale three-modal (text, image, voice) dataset, comprising 245,895 human-generated instances and 1,229,475 corresponding machine-generated instances. These machine-generated instances are produced using five current state-of-the-art (SOTA) generative models. All data instances are aligned across text, image, and voice modalities.
  • We introduce a variant of the proposed dataset specifically designed for model robustness testing. This version incorporates additional noise across all modalities, with three distinct types of noise applied to each modality.
  • We develop baseline detection models and evaluate their performance across all modalities. Our experimental results and analysis reveal that current SOTA triple-modality models struggle to accurately verify content originality and lack robustness against noise.
@@ 2. RU-AI Dataset
We introduce RU-AI, a new large-scale dataset that features multimodal real and human-generated content, accompanied by machine-generated content. It encompasses three modalities: real-world images, image-describing caption text, and spoken voices for captions, providing more than 245k real and machine-generated aligned instance pairs for each modality. More specifically, image and text instances are collected from COCO (Lin et al., 2014), Flickr8K (Young et al., 2014) and Places205 (Zhou et al., 2014), while voice instances are sourced from SLS Corpora  [^1] https://sls.csail.mit.edu/, which form the real and human-generated part of our dataset. The machine-generated portion is produced by five SOTA generative models specific to each modality. Figure 1 presents the data construction pipeline.
@@ 2.1. Data Construction
Voice. The voices for captions are cloud-sourced from a group of individuals, with each individual being assigned random captions  [^1] https://openai.com/index/chatgpt/. Some text-to-speech (TTS) models require reference voices for inference; therefore, we merge the spoken caption voices of each individual, identified by their unique user ID, into one-minute long voice files. For YourTTS (Casanova and et al., 2022), the machine-generated voices are synthesised by using the captions along with their corresponding one-minute human voices as references. For other models, the voices are synthesised directly from the caption text using default settings. All five models contribute equally to the generated voice data, and these voices are named and saved similarly to the images, with the original IDs and model names attached.
Data Augmentation. Based on the collected multimodal data, a noise-augmented variant is created towards robustness evaluation. For each modality, we introduce three different types of noise, with each type applied to 20% of the data. As a result, the final dataset consists of 60% for noise data and 40% clean data. For text, the noises add at the word level are: 1) randomly removing a single character from words; 2) randomly shuffling two characters within words; 3) randomly replacing a single character in words. These alterations are applied to 20% of the words in a sentence, and the augmented sentences adhere to the specified noise distribution. For images, three types of noise are added to randomly selected images through scikit-image (van der Walt et al., 2014), following the specified noise distribution: 1) Gaussian noise; 2) Salt and Pepper noise; 3) Poisson noise. For voices, three different noises are added with a 15dB Signal-to-Noise Ratio (SNR) through Audiomentations (Jordal et al., 2024), following the specified noise distribution: 1) Gaussian White noise; 2) Cafe noise; 3) Rain noise. More details can be found in the supplementary material.
@@ 2.2. Data Processing
We employ a regular expression method to remove LLM-generated greetings, such as ”Sure, here is the sentence:” and ”Certainly, here’s a rephrased version of your text:”, from the output to maintain the quality of the text data. Additionally, as generative models may produce empty sentences or blank images due to safety and alignment-related censorship, we remove these to ensure consistency across our dataset. To ensure cross-modality alignment, we perform a data-matching process that retains the minimal overlapping subset across all three modalities. Therefore a matched data instance is formed by aligning the image, text, and voice from different modality models corresponding to a specific ID. We implement this approach in our collected dataset, resulting in 245,895 aligned real-and-AI data pairs. As shown in Table 1, the final processed data for each modality includes 7,890 pairs from Flickr8K, 119,264 pairs from COCO, and 118,741 pairs from Places205.
@@
Text ChatGPT  [^1] https://openai.com/index/chatgpt/ 1,610 24,272 23,736
Llama-2-13B (Touvron et al., 2023) 1,518 23,121 23,488
@@
Mixtral-8x7B (Jiang and et al., 2024) 1,599 23,752 23,858
PaLM-2 (Anil and et al., 2023) 1,535 24,221 23,879
@@
AbsoluteReality-1.8  [^2] https://huggingface.co/Lykon/AbsoluteReality 1,628 23,898 23,780
@@
StableDiffusion-XL (Podell et al., 2024) 1,535 24,221 23,879
Voice YourTTS (Casanova and et al., 2022) 1,610 24,272 23,736
XTTS-2(Casanova et al., 2024) 1,518 23,121 23,488
@@ 2.3. Data Characteristics
Our dataset includes 245,895 pairs of real and machine-generated data from text, image, and voice modalities. Table 2 shows the average length of captions from different LLMs and their average similarity between the reference and generated text, indicated by average ROUGE-L and BERTScore. Table 3 shows the synthesised voice characteristics from different models, including average voice duration, average amplitude, and average spectral flatness. The average amplitude range refers to the average normalised audio scale, while the average spectral flatness is a measure of the sound’s noise-like level. Since all images in Flickr8K are unlabelled with object instances, we employ pre-trained YOLOv5 (Jocher and et al., 2021) to annotate the unlabelled images. The data characteristics are based on the original annotations and the YOLOv5 annotations. Our dataset can be divided into 20 different categories, with ”Person” being the largest category, and there are 1,040 images that do not belong to any category and are classified as ”Other”.
@@ 3.1. Implementation Details
We use two pre-trained SOTA multimodal encoder models with their large settings as the baseline multimodal embedding models: ImageBind (Girdhar et al., 2023) and LanguageBind (Zhu et al., 2024). ImageBind encodes all input data into 1,024-dimensional embedding sequences, while LanguageBind encodes them into 768-dimensional embedding sequences. Each multimodal embedding model is connected to a separate multilayer perceptron network, consisting of a single 256-dimensional hidden layer. The input layers of these networks match the dimensions of their respective embedding models, 1,024-dimensional for ImageBind and 768-dimensional for LanguageBind. These networks are used for the classification of the input data as either real or machine-generated one. Specifically, the output is calculated from probability vectors $\vec{z}=(z_{r},z_{m})$ using the softmax function $\sigma(\vec{z})_{i}$. The instance is classified as real if $z_{r}>z_{m}$; otherwise, it is classified as machine-generated. To analyse the effectiveness of different modalities in our data, we train and evaluate our classification model with different combinations of the data. The training samples are randomly selected from 80% of each data combination, while the remaining 20% of the samples are used as testing sets. We maintain the same size of data throughout the experiment and distribute the modalities equally. The performance is evaluated using Accuracy, Precision, and F1 metrics on both original and noise-augmented data. For the baseline implementation, we use PyTorch 1.13 with CUDA 11.8 on a server with a single Nvidia RTX 3090 GPU. All models were trained for 5 epochs, using the Adam optimizer with cross-entropy loss and a learning rate of 1e-4, with the embedding model frozen. We follow the embedding models’ original feature extraction techniques in all three modalities for consistency across datasets.
@@
Table 4. Results of the proposed classification model with ablation studies. ”LB” is LanguageBind, ”IB” is ImageBind, ”A” is Accuracy, ”P” is Precision, ”ALL” is all three modalities.
Model Modality Original Noise Augmentation
A@ALL P@ALL F1@ALL A@ALL P@ALL F1@ALL
LB + MLP Text 69.76 69.78 69.75 65.18 62.46 68.70
Image 76.49 85.95 72.93 62.98 86.95 45.32
Voice 68.89 64.09 73.43 68.36 63.66 73.07
Text+Image 77.09 76.91 77.16 64.63 71.76 57.82
Text+Voice 77.69 81.56 76.23 74.84 77.65 73.54
Image+Voice 82.26 78.89 83.23 78.20 91.34 74.12
All 84.20 84.20 82.67 80.02 78.02 80.74
IB + MLP Text 60.70 57.36 67.96 64.21 61.42 68.21
Image 67.75 71.03 65.03 65.54 63.51 68.04
Voice 65.77 63.61 68.28 66.93 63.39 70.87
Text+Image 71.48 75.18 69.23 68.58 72.41 65.73
Text+Voice 75.39 69.05 78.90 73.05 67.14 77.06
Image+Voice 81.25 82.84 80.78 79.54 75.99 80.88
All 80.63 79.19 81.10 76.42 72.97 78.11
3.2. Results
The classification results from our proposed model are presented in Table 4. To evaluate the effectiveness of each modality, we conduct ablation studies using single and mixed modalities during training. In general, baseline models that utilise LanguageBind as the embedding model outperform those incorporating ImageBind. The best F1 score achieved by LanguageBind MLP is 84.20, while the best F1 score from ImageBind MLP is 81.10. This may be attributed to LanguageBind’s enhanced cross-modality architecture, which binds all modalities through language, leveraging its rich semantics to provide more comprehensive information despite a smaller embedding size. For both baseline models, training with multi-modality data generally yields better performance compared to single-modality data. Since training with mixed-modality data introduces greater complexity compared to single-modality data, which could lead to better model convergence. We also observe that image and voice data tend to contribute more significantly to performance improvements. This may be caused by the richer features inherent in image and voice data compared to text. The performance trends are similar between the original data and the noise-augmented data across modalities. However, all baseline models exhibit decreased performance when evaluated on noise-augmented data, with an average drop of 2-4%. This decrease could be attributed to the added noise disrupting data features from both human and machine sources, which can subsequently lead to a degradation in model performance during evaluation.
Our baseline results from SOTA multimodal pre-trained models indicate that machine-generated content detection remains a challenging task, particularly for noise-augmented data. To mitigate the potential misuse of generative models, future work should focus on developing more robust and reliable detection methods.